OpenStudy (anonymous):

Can someone explain this for me step by step? Abby rows 10 km upstream and 10 km back in a total time of 3 hr. The speed of the river is 5 km/h. Find Abby’s speed in still water. @mark_o.

OpenStudy (anonymous):

You found me lol

OpenStudy (anonymous):

yeah im here now lol :D 

OpenStudy (anonymous):

she row from starting point A to B then from B to A that is time from A to B + time from B to A = 3 hours

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

time Tab + T ba = 3 (1) but distance d=velocity x time,. therefore t=d/v dAB dBA ------ + ------- =3 but distances dBA=dAB=10 km v v 10 10 ------- + ------ =3 v v but note that velocity of river is v=5 km/hr for upstream v= v+5 10 10 for down stream v= v-5, therefore ------- +------ = 3 v+5 v-5 you can algebra now ion finding v=?

OpenStudy (anonymous):

i mean you can do the algebra now from this point ... :D

OpenStudy (anonymous):

ummmm I am stuck

OpenStudy (anonymous):

I am sorry lol

OpenStudy (anonymous):

10/(v+5) + 10 /(v-5) =3 10(v-5) + 10(v+5) ------------------ =3 (v+5)(V-5)

OpenStudy (anonymous):

do multiplication and addition to simplify the numerator and the denominator..:D

OpenStudy (anonymous):

give me a sec

OpenStudy (anonymous):

wow i suck cause I am lost

OpenStudy (anonymous):